Mir Status Report 22 September 1997

Mir 24 / NASA 5 Status
Date: Monday, September 22, 1997
Mission Day: Mir 24 / NASA 5 MD 48/130

Mir 24 CDR Anatoly Solovyov
Mir 24 FE Pavel Vinogradov
NASA 5 FE Mike Foale

Mir Station System/Subsystem Activities

At 04:23 DMT this morning the motion control computer BTsVM-1 failed when two of three channels in the CPU failed. The gyrodines spun down and Elektron was shut down, but the ventilation system and thermal control loops remained active. The crew used Soyuz to initiate a spin, and within a few orbits the solar array current was 570 amps. The crew replaced Channel A with Channel C from the old computer, but it could not be determined which of the remaining two channels had failed, and there are no spares left onboard. However, the computer passed a self-test at 15:27 DMT, and the ground began reloading the software during the 18:21 DMT comm pass. The attitude reference was rebuilt during the DMT 23:01 comm pass, after which time thrusters-only attitude control was regained. Gyrodine spin-up began during the 00:28 pass.

The crew reported that they observed brown drops floating outside the vehicle around 10:00 DMT this morning, shortly after the Soyuz spinup. These drops seemed to emanate from the vicinity of Kvant-1 and moved in the direction of the other modules at about 20 deg. from the X axis. Mike reported they had a oily, shiny appearance, and that the previous crew had reported a similar sighting. Telemetry did not indicate any leaks.

Vozdukh continued to experience problems after being started on two vacuum valves (vacuum valve 2 was declared failed Saturday night). Further troubleshooting isolated the problem to the fan, which the crew replaced in the afternoon. This replacement did not solve the problem; the crew continued work, including looking for a new fan (there should apparently be two available onboard) until sleep. Work on Vozdukh will continue tomorrow.

No further maintenance was performed on the urine water recovery system SRV-U today.

Mir Payload Activities

Greenhouse Daily Ops (FE-1)
Daily Exercise (Treadmill) with recording on 'ZY'

OTHER ACTIVITIES
PFC for Mike

Data Management
On-Board
Nothing to report.

Network Status
Nothing to report.

Ground Facilities Status
Nothing to report.

Operations
Procedures/Timelines
Nothing received today.

Experiment Operations
No activities today due to the systems activities. Mike did report that both
Greenhouse and Beetle remain powered.

Experiment Hardware
Nothing to report.

Radiograms
MiSDE Checkout/Docking and MiSDE Docked phase activities radiograms were
completed and submitted for translation.

BTS checkout/card switch radiograms were completed and submitted for
translation.

Stowage/Transfer

MOST clarified to Mike over the weekend that he does not need to do anything with regards to the EVA safety tether. All work related to that item will be done during the EVA or after and Mike will not be involved. Any further questions regarding this item should be directed to the EVA group in Houston.

SMP/Medical

Mike participated in the second day of the physical training cycle with recording on the 'ZY'.

An audio only Private Family Conference (PFC) took place beginning at 19:58 DMT.

Environmental Parameters (Previous 24 hour period):
Parameter Min. Max. Nominal Range
Pressure (mmHg) Base Block 617 620 660-850
ppO2 (mmHg) 149 152 140-200
ppCO2 (mmHg) 4.8 7.3 up to 8.0
ppH20 (mmHg) 14.2 15.2 8-14
Humidity (percentage) 56 58 30-75
Temperature (xC) base block (BB) 25.2 26.6 18-28

Communications Summary
A/G with Consultants Group
19:58-20:08 PFC

23:10-23:12 Packing; BTS; MIM

00:42-00:46 Packing; Priroda drying ok

Packet --
The form 24 and one work file was submitted for uplink. Two files were received from Mike; one personal and one work (prepacking info).

Issues
None

Planning
MIR 24/NASA 5 MD 49/131
MIM Checkout

  • submit to reddit


Loading



Humans to Mars Summit 2019
Military Space Situational Awareness April 1 - April 2, 2019, London, UK
MilSatCom USA conference will return to Washington from the 26th – 27th June 2019
Space Access Conference 2019
Mission To Methonē by Les Johnson - Baen Books




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on September 22, 1997 8:40 PM.

Mir Status Report 20 September 1997 was the previous entry in this blog.

Mir Status Report 23 September 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.